I am in Algebra 2 currently and most things are pretty easy and I understand them well, but I'm having some trouble with parabolas, because I haven't worked on them since like 8th grade.
The equation is
y=2x^2+x
And I need it in vertex form.

Answer by lwsshak3(11628)

y=2x^2+x
And I need it in vertex form.
**
Standard form of equation for a parabola: y=A(x-h)^2+k, (h,k)=(x,y) coordinates of the vertex.
for given equation: y=2x^2+x
complete the square
y=2(x^2+x/2+1/16)-1/8
y=2(x+1/4)^2-1/8
vertex: (-1/4 -1/8)
This is a parabola that open up with vertex at (-1/4 -1/8)